Find two consecutive integers such that the product of the two integers is 12 more than the square of the smaller integer?
Answer by mananth(16946)

let the integers be x & x+1
x(x+1)=x^2-12
x^2+x=x^2-12
x=-12
The numbers are -12 & -13